# The Origin Symphony: Probing Baryogenesis with Gravitational Waves

### v1

@@ I. AD scalars: background dynamics and asymmetry generation
.5 Asymmetry dependence of GW production
It is worth exploring in more detail the connection between the asymmetry of the potential and the shape and amplitude of the resulting GW spectrum. Let us start from the symmetric case $A=0$, where the potential becomes
@@
Since the potential is radially symmetric, we can redefine the fields as $\phi_{\parallel}=\phi_{R}\cos\theta-\phi_{I}\sin\theta$ and $\phi_{\perp}=\phi_{R}\sin\theta+\phi_{I}\cos\theta$, where at the background level $\bar{\phi}_{\perp}=0$. This essentially makes the analysis follow our earlier work Cui et al. (2024), where we considered (among others) systems with two scalar fields with a quartic coupling, where one field is initially displaced from its minimum, and the other is (classically) at zero. This specific example is a combination of Models A and D of Ref. Cui et al. (2024). The equations of motion for the fluctuations in the $\phi_{\parallel}-\phi_{\perp}$ basis are
Figure S2: The spectra of the scalar field fluctuations along (left) and perpendicular (right) to the direction of the background trajectory in the case of a rotationally symmetric potential with $A=0$, $\lambda_{\Phi}=5\times 10^{-35}$ and $m_{\Phi}=0.1$ GeV.
@@
In principle, we can see two sources of parametric resonance: self-resonance of the ϕsubscriptitalic-$\phi_{\parallel}$ field due to the quartic term $3\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\parallel}$ and resonant amplification of the ϕsubscriptitalic-$\phi_{\perp}$ field through the coupling term $\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\perp}$.
@@
The background field trajectory is (by construction) along $\bar{\phi}_{\perp}=0$ and thus a straight line through the origin on the $\phi_{R}-\phi_{I}$ plane. Floquet analysis of a quartic field in a nonexpanding universe predicts that the resonance is stronger for $q=1$  ($\mu_{k}^{max}\sim 0.15$) than $q=3$ ($\mu_{k}^{max}\sim 0.04$Greene et al. (1997), where $q$ is defined as the coefficient before $\lambda_{\Phi}\bar{\phi}^{2}$ in Eqs. S33. Figure S2 shows the spectra of $\delta\phi_{\parallel},\delta\phi_{\perp}$ for different times, where we see that the $\delta\phi_{\perp}$ spectrum dominates, as expected from the standard analysis of preheating of quartic fields Greene et al. (1997). In the $\phi_{R}-\phi_{I}$ basis, the spectra will be identical for $\theta=\pi/4$, since each spectrum has an equal contribution from $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$. If the initial angle between the real and imaginary parts of $\Phi$ changes, the spectra of $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$ remain unchanged, whereas the spectra of $\phi_{R}$ and $\phi_{I}$ change, since the inner product between the basis vectors of the two bases change with the angle $\theta$. That being said, the resulting GW spectrum is unchanged, as expected, due to the rotational symmetry of the potential.
We now move to the case of $A\neq 0$. We parametrize the effect by $\epsilon_{\Phi}=A/m_{\Phi}^{2}$. As expected, small values of the angle do not significantly affect the background trajectory and the resonance structure. As seen in Fig. S3, for $\epsilon_{\phi}\leq 0.005$, the GW spectrum is independent of the initial angle $\theta_{\mathrm{in}}$ and actually indistinguishable from the symmetric case with $\epsilon_{\Phi}=0$. This means that for the asymmetry values that we use for our benchmark examples $\epsilon_{\Phi}<10^{-5}$, the resonance is “blind” with respect to the exact value of asymmetry in the potential.
@@
Figure S3: The resulting GW spectra for different values of the asymmetry parameter $A$ (color-coded according to the legend) and initial angle $\theta\equiv\arctan(\phi_{I}/\phi_{R})|_{\mathrm{init}}=\pi/6,\pi/4,\pi/3$ (solid, dashed and dotted respectively).
@@
That being said, it is worth pushing the asymmetry value higher in order to cover the entire possible parameter space. For $0.05\lesssim\epsilon_{\Phi}\lesssim 0.15$ we see an enhancement of the GW spectrum at large values of the frequency $f\gtrsim 100$ Hz, while the independence of $\theta_{\mathrm{in}}$ still holds. This is an intriguing observation, as it provides an additional motivation for GW detectors that can probe frequencies around $1$ kHz with a sensitivity similar to CE and ET Zhang et al. (2023); Chen et al. (2023). For $\epsilon_{\Phi}\simeq 0.25$ the UV enhancement of the GW spectrum remains, but we also see a difference between the initial angle choices between the real and imaginary parts of $\Phi$. This means that a possible detection of a spectrum with such a UV enhancement would suffer from a partial degeneracy between $\epsilon_{\phi}$ and $\theta_{\mathrm{in}}$. Since we focused on $\epsilon_{\Phi}\ll 1$, such a highly model-dependent calculation is outside the scope of the present work. Finally, for $\epsilon_{\Phi}\gtrsim 0.3$, parametric resonance is severely suppressed, rendering the GW signals below the detection threshold.

### v2

@@ Abstract
.5 Parametric resonance and asymmetry dependence of GW production
It is worth exploring in more detail the connection between the asymmetry of the potential and the shape and amplitude of the resulting GW spectrum. The growth of fluctuations and their backreaction breaks the homogeneous oscillation of $\Phi$, leading to scalar field inhomogeneities with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{b}\nabla_{j}\phi^{b}$ (where the repeated index b``b" indicates the number of the scalar components and a summation is implied). These time-dependent field inhomogeneities source (sub-Hubble) gravitational waves whose frequency depends on the characteristic energy scale involved $H_{c}$. The peak amplitude and frequency of GW spectra generated from sources peaked at some characteristic scale $k_{c}$ in momentum space can be estimated as Giblin and Thrane (2014)
$f_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.7\times 10^{10}\sqrt{\frac{H_{c}}{M_{\mathrm{Pl}}}}\frac{k_{c}}{H_{c}}\,{\mathrm{Hz}}\,$, (S31)
$\Omega_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.3\times 10^{-4}\,\alpha^{2}\,\beta\,w\left(\frac{k_{c}}{\sigma}\right)\left(\frac{H_{c}}{k_{c}}\right)^{2}\,$, (S32)
where $\alpha$ is a measure of the fraction of the energy in the GW source – here the scalar field inhomogeneities – relative to the total energy density of the Universe, and $\beta$ encodes the degree of anisotropy of the source, which must be determined from simulations. In typical scenarios $\beta=\mathcal{O}(0.01-0.1)$. The equation of the state of the Universe at the time of the production of GWs for a radiation-dominated Universe is $w=1/3$ and $\sigma$ is the width of the source in momentum space, which can be taken as $\sigma\sim k_{c}$ for peaked sources (see Ref. Giblin and Thrane (2014) for the derivation and e.g. Refs. Cui and Sfakianakis (2023); Cui et al. (2024) for further applications of Eqs (S31), (S32)).
Let us start from the symmetric case $A=0$, where the potential becomes
@@
Since the potential is radially symmetric, we can redefine the fields as $\phi_{\parallel}=\phi_{R}\cos\theta-\phi_{I}\sin\theta$ and $\phi_{\perp}=\phi_{R}\sin\theta+\phi_{I}\cos\theta$, where at the background level $\bar{\phi}_{\perp}=0$. This essentially follows our earlier work Cui et al. (2024), where we considered (among others) systems with two scalar fields featuring a quartic coupling, with one field initially displaced from its minimum and the other (classically) at zero. This specific example combines Models A and D from Ref. Cui et al. (2024). The equations of motion for the fluctuations in the $\phi_{\parallel}-\phi_{\perp}$ basis are
Figure S2: The spectra of the scalar field fluctuations along (left) and perpendicular (right) to the direction of the background trajectory in the case when the potential has a rotation symmetry with $A=0$, $\lambda_{\Phi}=5\times 10^{-35}$ and $m_{\Phi}=0.1$ GeV.
@@
In principle, we can see two sources of parametric resonance: self-resonance of the $\phi_{\parallel}$ field due to the quartic term $3\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\parallel}$ and resonant amplification of the $\phi_{\perp}$ field through the coupling term $\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\perp}$.
@@
The background field trajectory is (by construction) along $\bar{\phi}_{\perp}=0$ and thus a straight line through the origin on the $\phi_{R}-\phi_{I}$ plane. Floquet analysis of a quartic field in a nonexpanding universe predicts that the resonance is stronger for $q=1$  ($\mu_{k}^{max}\sim 0.15$) than $q=3$ ($\mu_{k}^{max}\sim 0.04$Greene et al. (1997), where $q$ is defined as the coefficient before $\lambda_{\Phi}\bar{\phi}^{2}$ in Eqs. S35. Figure S2 shows the spectra of $\delta\phi_{\parallel},\delta\phi_{\perp}$ for different times, where we see that the $\delta\phi_{\perp}$ spectrum dominates, as expected from the standard analysis of preheating of quartic fields Greene et al. (1997). In the $\phi_{R}-\phi_{I}$ basis, the spectra will be identical for $\theta=\pi/4$, since each spectrum has an equal contribution from $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$. If the initial angle between the real and imaginary parts of $\Phi$ changes, the spectra of $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$ remain unchanged, whereas the spectra of $\phi_{R}$ and $\phi_{I}$ change, since the inner product between the basis vectors of the two bases changes with the angle $\theta$. That being said, the resulting GW spectrum is unchanged, as expected, due to the rotational symmetry of the potential.
We now move to the case of $A\neq 0$. We parametrize the effect by $\epsilon_{\Phi}=A/m_{\Phi}^{2}$. Let us first discuss the background trajectory. For $\epsilon_{\Phi}=0$ it is a straight line segment through the origin, with a length of ${\cal O}(\phi_{\mathrm{in}})$ which will shrink in time due to red-shifting. For $0<\epsilon_{\Phi}\ll 1$, the trajectory during each period will resemble an ellipse. Using Eqs. (S15)-(S16)) and Taylor-expanding for $\epsilon_{\Phi}\ll 1$, we can see that the minor axis of the (approximate) ellipse scales as ${\cal O}(\phi_{\mathrm{in}}\cdot A/m_{\phi}^{2})$, whereas the major axis is simply ${\cal O}(\phi_{\mathrm{in}})$. The ellipticity is then $\sim(A/m_{\phi}^{2})\ll 1$ for all cases of interest shown in the main text. For small values of $\epsilon_{\Phi}$, the background trajectory will be an ellipse with a very small ellipticity, and it will therefore behave almost like a straight line segment. We thus expect the parametric resonance results to be unaffected by the exact value of $\epsilon_{\Phi}\ll 1$. As seen in Fig. S3, for $\epsilon_{\phi}\leq 0.005$, the GW spectrum is independent of the initial angle $\theta_{\mathrm{in}}$ and indistinguishable from the symmetric case with $\epsilon_{\Phi}=0$. This means that for the asymmetry values that we use for our benchmark examples $\epsilon_{\Phi}<10^{-5}$, the resonance is “blind” with respect to the exact value of asymmetry in the potential.
@@
Figure S3: The resulting GW spectra for different values of the asymmetry parameter AA (color-coded according to the legend) and initial angle $\theta\equiv\arctan(\phi_{I}/\phi_{R})|_{\mathrm{init}}=\pi/6,\pi/4,\pi/3$ (solid, dashed and dotted respectively).
@@
That being said, it is worth pushing the asymmetry value higher in order to cover the entire possible parameter space. For $0.05\lesssim\epsilon_{\Phi}\lesssim 0.15$, we see an enhancement of the GW spectrum at large values of the frequency $f\gtrsim 100$ Hz, while the independence of $\theta_{\mathrm{in}}$ still holds. This is an intriguing observation, as it provides an additional motivation for GW detectors that can probe frequencies around $1$ kHz with a sensitivity similar to CE and ET Zhang et al. (2023); Chen et al. (2023). For $\epsilon_{\Phi}\simeq 0.25$, the UV enhancement of the GW spectrum remains, but we also see a difference between the initial angle choices between the real and imaginary parts of $\Phi$. This means that a possible detection of a spectrum with such a UV enhancement would suffer from a partial degeneracy between $\epsilon_{\phi}$ and $\theta_{\mathrm{in}}$. Since we focused on $\epsilon_{\Phi}\ll 1$, such a highly model-dependent calculation is outside the scope of the present work. Finally, for $\epsilon_{\Phi}\gtrsim 0.3$, parametric resonance is severely suppressed, rendering the GW signals below the detection threshold.
Before concluding, we must note that since $\Phi$ is coupled to SM fermions in order to transfer the asymmetry to baryons, the effect of parametric excitation of these fermionic states should be evaluated. Nevertheless, in the context of preheating, it was found that during scalar field oscillations, the energy transfer to bosons is typically faster and larger than to fermions due to Pauli blocking Greene and Kofman (1999); Giudice et al. (1999); Greene and Kofman (2000); Sorbo (2001); Adshead and Sfakianakis (2015). Therefore, we expect our resonance analysis to remain robust in the presence of fermion couplings.